OpenStudy (anonymous):

Will give medals! and become fan, please help~ use de moivre's theorem to write the complex number in trigonometric form. ((sqrt3)+i)^3

Parth (parthkohli):

Have you learned any methods to do this?

OpenStudy (anonymous):

A. 2(cos(5pi/3) + isin(5pi/3)) B. 2(cos(2pi/3) + isin(2pi/3)) c. 8(cos(pi) + isin(pi)) D. 8(cos(pi/2) + isin(pi/2))

Parth (parthkohli):

OK. Do you know how to do this?

OpenStudy (anonymous):

and just Zn=r^n(cos(ntheta) + isin(ntheta), but other than that i have no clue

Parth (parthkohli):

Alright.\[\left(\sqrt 3 + i\right)^3\]\[= \left(2 \left(\frac{\sqrt{3}}{2} + \frac{1}{2} i\right)\right)^3\]

OpenStudy (anonymous):

okay, im following thus far..

Parth (parthkohli):

Is the expression\[\frac{\sqrt{3}}{2} + \frac{1}{2}i\]in the form of \(\cos\theta + i\sin\theta\)? Do you know which \(\theta\) in particular?

OpenStudy (anonymous):

honestly, im not sure

Parth (parthkohli):

\[\cos\theta = \dfrac{\sqrt{3}}{2}\]and\[\sin\theta = \dfrac{1}{2}\]Do you happen to know an angle that satisfies both of these?

OpenStudy (anonymous):

pi/6?

OpenStudy (anonymous):

write ((sqrt3)+i)^3 in trigonometric form, how can i accomplish that?

Parth (parthkohli):

That's correct.\[\left(2 \left(\frac{\sqrt{3}}{2} + \frac{1}{2} i\right)\right)^3\]\[= 8 \times \left(\cos\left(\frac{\pi}{6}\right)+ i\sin\left(\frac{\pi}{6}\right)\right)^3 \]Does that remind you of De Moivre's Theorem?

OpenStudy (anonymous):

yes, thats very familiar!

OpenStudy (anonymous):

what does it mean since pi/6 isnt in one of my answer options?

Parth (parthkohli):

Let me restate De Moivre's Theorem:\[\left(\cos\theta + i\sin\theta\right)^n = \cos(n\theta) + i\sin(n\theta)\]As you can see, \(n = 3\).

OpenStudy (anonymous):

because the nth root of a unit for any value is n?

Parth (parthkohli):

I don't get what you're asking, but...\[= 8 \times \left(\cos\left(\frac{\pi}{6}\right)+ i\sin\left(\frac{\pi}{6}\right)\right)^3\]\[= 8 \times \left(\cos\left(\frac{\pi}{6}\times 3\right) + i\sin\left(\frac{\pi}{6}\times 3\right)\right)\]

OpenStudy (anonymous):

my original answer would have been 8(cos(pi/2) + isin(pi/2))

OpenStudy (anonymous):

and yeah, thats okay! XD

Parth (parthkohli):

=)

OpenStudy (anonymous):

thank you!

Parth (parthkohli):

You're welcome :)